OpenStudy (anonymous):

solve x^3 is less than 25x

OpenStudy (anonymous):

x^3 < 25 x x^3 - 25 x < 0 x (x^2 - 25) < 0 for x > = 5, x^2 - 25 > = 0 and x too will be > = 0, so x (x^2 - 25) > = 0 for x = < -5, x^2 - 25 > = 0 and x < = 0, so x (x^2 - 25) < = 0. Your answer: x < - 5.

OpenStudy (anonymous):

Not one of my possible answers.

OpenStudy (anonymous):

Really? What's the answer in the book?

OpenStudy (anonymous):

1)-5<x<0 or x>5 ,/ 2) -5<x<5 /3) x<-5 orx>5 /4) x<-5or 0<x<5

OpenStudy (anonymous):

Or x >0 and x <5.

OpenStudy (anonymous):

4) x<-5or 0<x<5

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You're welcome. :)